John Q. is paid bi-weekly and earns $1,125 per paycheck. If the payment system at John's work changes to semi-monthly , how much would each of his paychecks now be?
Answer by RAY100(1637)

This problem revolves around the fact that bi- weekly (with 52 weeks / yr) is not the same as bi monthly(with 12 months / yr)
.
Yearly pay is, $1,125 / 2 weeks * 52 weeks / yr = $ 29,250
.
Bi monthly pay is ,,,$29,250 / yr * 1 yr/12 months * 1month/2 bi-monthly = $ 1218.75 bimonthly.
.
.
check, 1125 * 26 = $29,250,,,,,1218.75 *24 = $ 29.250,,,,,ok
.